A driver travels northbound on a highway at a speed of 25.0m/s. A police car, travelling southbound at a speed of 40.0m/s, approaches with its siren producing sound at a frequency of 2500Hz

(a) What frequency does the driver observe as the police car approaches?

(b) What frequency does the driver detect after the police car passes him?


1
Expert's answer
2021-02-22T10:28:44-0500

(a)

"f_o=f_s(\dfrac{v+v_o}{v-v_s}),""f_o=2500\ Hz\cdot(\dfrac{343\ \dfrac{m}{s}+25\ \dfrac{m}{s}}{343\ \dfrac{m}{s}-40\ \dfrac{m}{s}})=3036.3\ Hz."


(b)

"f_o=f_s(\dfrac{v-v_o}{v+v_s}),""f_o=2500\ Hz\cdot(\dfrac{343\ \dfrac{m}{s}-25\ \dfrac{m}{s}}{343\ \dfrac{m}{s}+40\ \dfrac{m}{s}})=2075.7\ Hz."